# C.I.T. Campus, Taramani, Chennai-600113 Tamil Nadu, India

### v1

@@ 4.1 The Frame-dragging
$\begin{split}g^{00}=&\frac{1}{\Delta}\left(r^{2}+\alpha^{2}+\frac{r_{g}r-Q^{2}}{\Sigma^{2}}\alpha^{2}\sin^{2}\theta\right),\\ g^{03}=&g^{30}=\frac{r_{g}r-Q^{2}}{\Delta\Sigma^{2}}\alpha,\\ g^{33}=&-\frac{\Sigma^{2}-(r_{g}r-Q^{2})}{\Delta\Sigma^{2}\sin^{2}\theta},\\ \end{split}$ (29)
@@
Thus, the frame-dragging of KNBH is obtained as {strip}
$\begin{split}\frac{d\phi}{cdt}=&\frac{\left[(r_{g}r-Q^{2})\alpha\sin^{2}\theta\frac{E_{0}}{c}\right]+\left[\Sigma^{2}-(r_{g}r-Q^{2})\right]L}{\left[\Sigma^{2}\left(r^{2}+\alpha^{2}+\frac{r_{g}r-Q^{2}}{\Sigma^{2}}\alpha^{2}\sin^{2}\theta\right)\frac{E_{0}}{c}-(r_{g}r-Q^{2})\alpha L\right]\sin^{2}\theta}\\ =&\frac{\left[(r_{g}r-Q^{2})\alpha\sin^{2}\theta\right]+\left[\Sigma^{2}-(r_{g}r-Q^{2})\right]b}{\left[\Sigma^{2}\left(r^{2}+\alpha^{2}+\frac{r_{g}r-Q^{2}}{\Sigma^{2}}\alpha^{2}\sin^{2}\theta\right)-(r_{g}r-Q^{2})\alpha b\right]\sin^{2}\theta}.\end{split}$ (32)
This is the general expression of the frame dragging of KNBH in the non-equatorial plane in a strong field limit. Here, it should be noted that in the propagation of a light ray (or a photon-like particle) parallel to the x-axis, i.e. $\phi$ is restricted to ±$\pm\pi/2$ (at the location where the impact parameter is measured), the momentum $(p)$ and the conserved energy $(E_{0})$ can be expressed by relation $E_{0}=pc$ and the angular momentum $L$ as $L=pb$, where impact parameter $=\sqrt{y^{2}+z^{2}}=r\sqrt{\sin^{2}\theta\sin^{2}\phi+\cos^{2}\theta}=r\sqrt{\sin^{2}\theta+\cos^{2}\theta}=r=b$ ($b$ is the physical radius of KNBH).
Now, under the far-field approximation the frame dragging of KNBH in the non-equatorial plane becomes,
$\frac{d\phi}{cdt}=\frac{\left[(r_{g}r-Q^{2})\alpha\sin^{2}\theta\right]+\left[r^{2}-(r_{g}r-Q^{2})\right]b}{\left[r^{4}-(r_{g}r-Q^{2})\alpha b\right]\sin^{2}\theta}.$ (33)
Furthermore, if we set $\theta=\frac{\pi}{2}$ in Eq.(33), the frame dragging for KNBH can be obtained in the equatorial plane as,
$\frac{d\phi}{cdt}=\frac{\left[(r_{g}r-Q^{2})\alpha\right]+\left[r^{2}-(r_{g}r-Q^{2})\right]b}{\left[r^{4}-(r_{g}r-Q^{2})\alpha b\right]}.$ (34)
By substituting the value of $\frac{d\phi}{cdt}$ from Eq.(33) into Eq.(24) or Eq. (25) the refractive index at an arbitrary point on the non-equatorial plane in the KNBH field is {strip}
$\begin{split}n(r,\theta,\alpha,Q)=&\left\{1-\frac{r_{g}}{r}+\frac{Q^{2}}{r^{2}}\right\}^{-1}\left[1+\frac{2(r_{g}r-Q^{2})\alpha}{r^{2}-r_{g}r+Q^{2}}\times\frac{\left[(r_{g}r-Q^{2})\alpha\sin^{2}\theta\right]+\left[r^{2}-(r_{g}r-Q^{2})\right]b}{\left[r^{4}-(r_{g}r-Q^{2})\alpha b\right]}\right]^{-1/2},\end{split}$ (35)
or,
$\begin{split}n(x,\theta,\alpha,Q)=&\left(\frac{x}{x-1}\right)\left\{1+\frac{q^{2}}{x(x-1)}\right\}^{-1}\left[1+\frac{2(x-q^{2})u}{x^{2}-x+q^{2}}\times\frac{\left[(x-q^{2})u\sin^{2}\theta\right]+\left[x^{2}-(x-q^{2})\right]v}{\left[x^{4}-(x-q^{2})uv\right]}\right]^{-1/2}\\ =&n_{0}(x)(1+C_{x})^{-1}(1+2S_{x})^{-\frac{1}{2}},\end{split}$ (36)
@@
The Eq.(35) or (36) is the general expression of the refractive index on the non-equatorial plane in the field of rotating charged body.
Now, it is possible to show that for all $r>>r_{g}$ and $r>>Q$, we must have $x>>1$ and $x>>q$. As $x>>1$, we can approximate $(x-1)\approx x$ and then we can finally show that $C_{x}<<1$. There is another possibility that, when $r>>r_{g}$ and $r>>\alpha$, we must have $x>>1$ and $x>>u$. Also, since $\alpha<b$, we must have $u<v$ and finally $S_{x}<<1$. So, from the assumption of $r>>r_{g}$, $r>>Q$ and $r>>\alpha$ that results $C_{x}<<1$ and $S_{x}<<1$, we can say $n<n_{0}$ i.e., the refractive index in the field of rotating charged body is less than in the field of static body.
If we set $\alpha=0$, the gravitational mass becomes charged and the above Eq.(35) or (36) goes to that for RN geometry roy2017deflection . If we set only the charge parameter $(q)$ as zero, then the refractive index in the equatorial plane ($\theta=\frac{\pi}{2}$) becomes that of the Kerr geometry roy2015trajectory . But if both charge and rotation parameters vanish, the central gravitational mass is static and uncharged and in that case, the above expression of the refractive index goes over to that of Schwarzschild sen2010more .
@@
Now, by substituting $\theta=\frac{\pi}{2}$ in Eq.(35) or (36), the refractive index due to KNBH can be obtained in the equatorial plane as
@@
Fig. 9 visualizes the variation of the refractive index with respect to the three key parameters: charge, rotation, and polar angle simultaneously. It serves as a comprehensive graphical summary of the preceding plots, enabling a unified understanding of how these parameters collectively influence the optical behavior of the KNBH. In the prograde case, the refractive index exhibits an increasing trend with the rotation parameter and a decreasing trend with charge. The effect is most pronounced at equatorial plane $(\theta=\pi/2)$, where the refractive index attains its maximum, indicating a stronger deflection. Conversely, in the retrograde case, the refractive index is generally reduced. It decreases with increasing rotation and increases with charge, showing inverse behavior compared to the prograde scenario. Additionally, the minimum refractive index typically occurs near the polar regions $(\theta=0,\pi)$. These plots collectively highlight the significant and nontrivial interplay between charge, rotation, and angular position in shaping the refractive index profile.
@@ 5 Deflection due to Kerr-Newman Black Hole
In the present case, light approaches from asymptotic infinity, parallel to x-axis ($r=-\infty$ or $x=-\infty$) toward a charged rotating gravitational mass at the origin, defined by the Schwarzschild radius $r_{g}$, rotation parameter $\alpha$, and associated charge $Q$. The beam reaches $r=\infty$ or $x=\infty$ after a specified amount of deflection ($\Delta\psi$). The closest approach distance to the incoming ray (known as the impact parameter) is the perpendicular distance between the center of the KNBH and the ray. Now, in our consideration, all the rays are cutting the y-z plane perpendicularly. On the y-z plane, for all the points, x-coordinate is zero and when the ray cuts the plane, $\phi=-\pi/2$ or $+\pi/2$. So, the perpendicular distance between the ray and the center of the KNBH is $\sqrt{y^{2}+z^{2}}=r=b$ (b is the physical radius of the KNBH), shown in Fig. 10. Now, we change the variable to $x=\frac{r}{r_{g}}$ so that $dr=r_{g}dx$ and the corresponding limit changes from $x=v$ to $x=\infty$, as the limit of $r$ changes from $r=b$ to $r=\infty$ in Eq.(41)
$\begin{split}\Delta\psi&=2\int_{v}^{\infty}\frac{dx}{x\sqrt{(\frac{n(x).x}{n(v).v})^{2}-1}}-\pi=2I-\pi,\end{split}$ (42)
where
$\begin{split}I&=\int_{v}^{\infty}\frac{dx}{x\sqrt{(\frac{n(x).x}{n(v).v})^{2}-1}}\\ &=n(v).v\int_{v}^{\infty}\frac{dx}{x\sqrt{(n(x).x)^{2}-(n(v).v)^{2}}}.\end{split}$ (43)
@@
Substituting the value of the refractive index from Eq.(36), we rewrite the above Eq. (43) of $I$ as
{strip}
$\begin{split}I&=D_{KN}\int_{v}^{\infty}\frac{dx}{x\sqrt{[n_{0}(x).x(1+C_{x})^{-1}(1+2S_{x})^{-\frac{1}{2}}]^{2}-D^{2}_{KN}}}\\ &=D_{KN}\int_{v}^{\infty}\frac{dx}{x\sqrt{n^{2}_{0}(x).x^{2}(1+C_{x})^{-2}(1+2S_{x})^{-1}-D^{2}_{KN}}}\\ &=D_{KN}\int_{v}^{\infty}\frac{dx}{x\sqrt{n^{2}_{0}(x).x^{2}-D^{2}_{0}+n^{2}_{0}(x).x^{2}(1+C_{x})^{-2}(1+2S_{x})^{-1}-n^{2}_{0}(x).x^{2}+D^{2}_{0}-D^{2}_{KN}}}\\ &=D_{KN}\int_{v}^{\infty}\frac{dx}{x\sqrt{n^{2}_{0}(x).x^{2}-D^{2}_{0}}}\left[1+\frac{n^{2}_{0}(x).x^{2}\{(1+C_{x})^{-2}(1+2S_{x})^{-1}-1\}+D^{2}_{0}-D^{2}_{KN}}{n^{2}_{0}(x).x^{2}-D^{2}_{0}}\right]^{-\frac{1}{2}}\\ &=D_{KN}\int_{v}^{\infty}\frac{dx}{x\sqrt{n^{2}_{0}(x).x^{2}-D^{2}_{0}}}[1+P(x)]^{-\frac{1}{2}},\end{split}$ (44)
with $D_{0}=n_{0}(v).v$, corresponding to SBH deflection, $n(v).v=D_{KN}=D_{0}(1+C_{v})^{-2}(1+2S_{v})^{-1}$ and also we have denoted
$P(x)=\frac{n^{2}_{0}(x).x^{2}\{(1+C_{x})^{-2}(1+2S_{x})^{-1}-1\}+D^{2}_{0}-D^{2}_{KN}}{n^{2}_{0}(x).x^{2}-D^{2}_{0}}.$ (45)
Eq. (42) may be used to obtain the exact deflection of light on the non-equatorial plane of a KNBH as
$\begin{split}\Delta\psi&=2I-\pi\\ &=2D_{KN}\int_{v}^{\infty}\frac{dx}{x\sqrt{n^{2}_{0}(x).x^{2}-D^{2}_{0}}}[1+P(x)]^{-\frac{1}{2}}-\pi.\end{split}$ (46)
The above deflection angle can be expanded as a series
$\begin{split}\Delta\psi=&2D_{KN}\int_{v}^{\infty}\frac{dx}{x\sqrt{n^{2}_{0}(x).x^{2}-D^{2}_{0}}}\left[1-\frac{1}{2}P(x)+\frac{3}{8}P^{2}(x)-\frac{5}{16}P^{3}(x)+.................\right]-\pi.\end{split}$ (47)
Also, the deflection of light ray in the equatorial plane can be obtained by considering the refractive index expressed by Eq. (39).
@@
Figs 11-13, collectively illustrate the variation of the deflection angle as a function of the impact parameter, charge parameter, rotation parameter and polar angle for KNBH and other BH solutions in GR. Fig. 11 shows that the deflection angle decreases monotonically with increasing impact parameter. Among the configurations, the deflection is most pronounced for the KBH in the prograde case and is least for the KNBH retrograde case. The physical significance of this result is that the deflection angle is largest for the KBH in the prograde case due to frame dragging aiding the photon’s approach, increasing spacetime curvature. Conversely, it is smallest for the KNBH retrograde case, where opposing frame dragging and charge both reduce the effective bending. Furthermore, the deflection angle for KNBH attains its maximum value on the equatorial plane for prograde motion and its minimum in the retrograde configuration. A comparative analysis with other BHs reveals the ordering of the deflection strength as: KBH prograde $>$ SBH $>$ KBH retrograde $>$ KNBH prograde $>$ RNBH $>$ KNBH retrograde. Fig. 12 depicts the variation of the deflection angle with the charge parameter. Since the KBH and SBH are independent of charge, their deflection angles remain constant and are represented by horizontal lines. From the graphical representation, it is once again evident that the KBH prograde configuration exhibits the most dominant deflection, while the KNBH retrograde case shows the least. For the KNBH alone, the deflection is more pronounced in the prograde configuration compared to the retrograde one, and it also exceeds that of the RNBH. Further, the maximum deflection is observed for the prograde motion at the equatorial plane and the minimum for the retrograde motion in the same plane. Fig. 13 presents the dependence of the deflection angle on the rotation parameter. In this case, the RNBH and SBH yield constant deflection angles due to the absence of rotational dependence, appearing as horizontal lines. For the KNBH, the deflection angle increases with rotation in the prograde case and decreases in the retrograde case. In addition, inducing charge parameters lead to a reduction in the deflection angle for both cases. Comparatively, the KBH (uncharged rotating BH) exhibits a stronger deflection than the KNBH, underscoring the diminishing influence of charge on the deflection angle. The dependency of polar angle on deflection angle remains consistent with previous cases, having reached a maximum on the equatorial plane and decreasing toward the polar plane.
@@ 6 Comparison with Kerr-Sen Black Hole
7 Summary and Conclusions
@@
These consistencies demonstrate that the material medi-um approach provides a robust and reliable framework for analyzing light deflection in rotating charged BH spacetimes.
@@ Acknowledgements.
Authors, SR and PG express sincere and deep gratitude to the Department of Physics, NITA, for providing the necessary research environment to complete this work. The author, SK, sincerely acknowledges IMSc for providing exceptional research facilities and a conducive environment that facilitated his work as an Institute Postdoctoral Fellow. One of the authors, HN, would like to thank IUCAA, Pune, for the support under its associateship program, where a part of this work was done. The author HN also acknowledges the financial support provided by the Science and Engineering Research Board (SERB), New Delhi, through grant number CRG/2023/008980.

### v2

@@ 4.1 The Frame-dragging
[…]
@@
The Eq.(LABEL:Eq_ref_r_FD) or (LABEL:Eq_ref_x_FD) is the general expression of the refractive index on the non-equatorial plane in the field of rotating charged body. Now, it is possible to show that for all $r>>r_{g}$ and $r>>Q$, we must have $x>>1$ and $x>>q$. As $x>>1$, we can approximate $(x-1)\approx x$ and then we can finally show that $C_{x}<<1$. There is another possibility that, when $r>>r_{g}$ and $r>>\alpha$, we must have $x>>1$ and $x>>u$. Also, since $\alpha<b$, we must have $u<v$ and finally $S_{x}<<1$. So, from the assumption of $r>>r_{g}$, $r>>Q$ and $r>>\alpha$ that results $C_{x}<<1$ and $S_{x}<<1$, we can say $n<n_{0}$ i.e., the refractive index in the field of rotating charged body is less than in the field of static body. If we set $\alpha=0$, the gravitational mass becomes charged and the above Eq.(LABEL:Eq_ref_r_FD) or (LABEL:Eq_ref_x_FD) goes to that for RN geometry roy2017deflection . If we set only the charge parameter $(q)$ as zero, then the refractive index in the equatorial plane ($\theta=\frac{\pi}{2}$) becomes that of the Kerr geometry roy2015trajectory . But if both charge and rotation parameters vanish, the central gravitational mass is static and uncharged and in that case, the above expression of the refractive index goes over to that of Schwarzschild sen2010more .
@@
Now, by substituting $\theta=\frac{\pi}{2}$ in Eq.(LABEL:Eq_ref_r_FD) or (LABEL:Eq_ref_x_FD), the refractive index due to KNBH can be obtained in the equatorial plane as
@@
Fig. 9 visualizes the variation of the refractive index with respect to the three key parameters: charge, rotation, and polar angle simultaneously. It serves as a comprehensive graphical summary of the preceding plots, enabling a unified understanding of how these parameters collectively influence the optical behavior of the KNBH. In the prograde case, the refractive index exhibits an increasing trend with the rotation parameter and a decreasing trend with charge. The effect is most pronounced at equatorial plane $(\theta=\pi/2)$, where the refractive index attains its maximum, indicating a stronger deflection. Conversely, in the retrograde case, the refractive index is generally reduced. It decreases with increasing rotation and increases with charge, showing inverse behavior compared to the prograde scenario. Additionally, the minimum refractive index typically occurs near the polar regions $(\theta=0,\pi)$.
These plots collectively highlight the significant and nontrivial interplay between charge, rotation, and angular position in shaping the refractive index profile.
@@ 5 Deflection due to Kerr-Newman Black Hole
In the present case, light approaches from asymptotic infinity, parallel to x-axis ($r=$-$\infty$ or $x=$-$\infty$) toward a charged rotating gravitational mass at the origin, defined by the Schwarzschild radius $r_{g}$, rotation parameter $\alpha$, and associated charge $Q$. The beam reaches $r=\infty$ or $x=\infty$ after a specified amount of deflection ($\Delta\psi$). The closest approach distance to the incoming ray (known as the impact parameter) is the perpendicular distance between the center of the KNBH and the ray. Now, in our consideration, all the rays are cutting the y-z plane perpendicularly. On the y-z plane, for all the points, x-coordinate is zero and when the ray cuts the plane, $\phi=-\pi/2$ or $\pi/2$. So, the perpendicular distance between the ray and the center of the KNBH is $\sqrt{y^{2}+z^{2}}=r=b$ (b is the physical radius of the KNBH), shown in Fig. 10.
@@
Figs 11-13, collectively illustrate the variation of the deflection angle as a function of the impact parameter, charge parameter, rotation parameter and polar angle for KNBH and other BH solutions in GR. Fig. 11 shows that the deflection angle decreases monotonically with increasing impact parameter. Among the configurations, the deflection is most pronounced for the KBH in the prograde case and is least for the KNBH retrograde case. The physical significance of this result is that the deflection angle is largest for the KBH in the prograde case due to frame dragging aiding the photon’s approach, increasing spacetime curvature. Conversely, it is smallest for the KNBH retrograde case, where opposing frame dragging and charge both reduce the effective bending. Furthermore, the deflection angle for KNBH attains its maximum value on the equatorial plane for prograde motion and its minimum in the retrograde configuration. A comparative analysis with other BHs reveals the ordering of the deflection strength as: KBH prograde >> SBH >> KBH retrograde >> KNBH prograde >> RNBH >> KNBH retrograde. Fig. 12 depicts the variation of the deflection angle with the charge parameter. Since the KBH and SBH are independent of charge, their deflection angles remain constant and are represented by horizontal lines. From the graphical representation, it is once again evident that the KBH prograde configuration exhibits the most dominant deflection, while the KNBH retrograde case shows the least. For the KNBH alone, the deflection is more pronounced in the prograde configuration compared to the retrograde one, and it also exceeds that of the RNBH. Further, the maximum deflection is observed for the prograde motion at the equatorial plane and the minimum for the retrograde motion in the same plane. Fig. 13 presents the dependence of the deflection angle on the rotation parameter. In this case, the RNBH and SBH yield constant deflection angles due to the absence of rotational dependence, appearing as horizontal lines. For the KNBH, the deflection angle increases with rotation in the prograde case and decreases in the retrograde case. In addition, inducing charge parameters lead to a reduction in the deflection angle for both cases. Comparatively, the KBH (uncharged rotating BH) exhibits a stronger deflection than the KNBH, underscoring the diminishing influence of charge on the deflection angle. The dependency of polar angle on deflection angle remains consistent with previous cases, having reached a maximum on the equatorial plane and decreasing toward the polar plane.
@@ 6 Comparison with Kerr-Sen Black Hole
7 Summary, Conclusions and Future directions
@@
  • $\bullet$
    Finally, we compare our findings with those of Wang et al. Wang:2022ouq , who obtained KNBH geodesics in terms of elliptic integrals and Jacobian elliptic functions for both null and timelike trajectories, using the Mino time formalism with specified initial conditions. Their analysis reveals boomerang-like spherical orbits arising from frame dragging and black hole charge. Our results exhibit similar qualitative trends, where the deflection decreases with increasing charge and impact parameter, increases with spin in prograde motion, and decreases in retrograde motion, consistent with their frame-dragging interpretation. The angular dependence we obtain, strongest in the equatorial plane and gradually diminishing toward the poles, also aligns with their conclusions. However, as the two studies employ different formalisms, a direct quantitative comparison is not possible and the agreement should be regarded as qualitative.
Future Directions
  • $\bullet$
    It is well established that extremal BHs exhibit distinctive features in their photon sphere structure and light deflection behavior. In particular, as demonstrated by Chen et al. Chen:2024oyv , the extremal KNBH permits a stable photon orbit exactly at the horizon due to the presence of a degenerate root in the radial potential. This leads to a qualitatively different behavior in the deflection angle compared to subextremal configurations, including strong deviations near the critical impact parameter. These properties are unique to the extremal limit and influence photon trajectories in ways not observed in non-extremal BHs. Although the present study does not consider the extremal regime, it would be interesting in future work to investigate such cases using the material medium approach.
@@ Acknowledgements.
We sincerely thank the anonymous referee for their careful review and constructive comments, which have improved the scientific content and overall quality of this paper. Authors, SR and PG express sincere and deep gratitude to the Department of Physics, NITA, for providing the necessary research environment to complete this work. The author, SK, sincerely acknowledges IMSc for providing exceptional research facilities and a conducive environment that facilitated his work as an Institute Postdoctoral Fellow. One of the authors, HN, would like to thank IUCAA, Pune, for the support under its associateship program, where a part of this work was done. The author HN also acknowledges the financial support provided by the Science and Engineering Research Board (SERB), New Delhi, through grant number CRG/2023/008980.